SE Homework 3 - Excel Sign in - o X Share File Home Insert Page Layout Formulas Data Review View Help Tell me what you want tCan someone help me solve 1 and 2?
Answer #1
x P(x) xP(x) x2P(x)
0 0.0323 0.0000 0.0000
9 0.0048 0.0435 0.3919
74 0.0006 0.0477 3.5329
304 0.0002 0.0490 14.9058
1279 0.0001 0.0825 105.5381
2699 0.0000 0.0871 234.9871
-1 0.9620 -0.9620 0.9620
total -0.6521 360.3179
E(x) =μ= ΣxP(x) = -0.6521
E(x2) = Σx2P(x) = 360.3179
Var(x)=σ2 = E(x2)-(E(x))2= 359.8927
std deviation=         σ= √σ2 = 18.9708

1)

expected profit = -0.6521

2)

standard deviation =18.9708

3)

here this is binomial distribution with parameter n=11 and p=1/5

probability = P(3<=X<=6)= x=x1x2     (nCx)px(1−p)(n-x)    = 0.3806